# Spin-orbit coupling in symmetric and mixed spin-symmetry

### v1

@@ Abstract
Synthetically spin-orbit coupling in cold atoms couples the pseudo-spin and spatial degrees of freedom, and therefore the inherent spin symmetry of the system plays an important role. In systems of two pseudo-spin degrees, two particles configure symmetric states and anti-symmetric states, but the spin symmetry can be mixed for more particles. We study the role of mixed spin symmetry in the presence of spin-orbit coupling and consider the system of three bosons with two hyper-fine states trapped in a harmonic potential. We investigate the ground state and the energy spectrum by implementing exact diagonalization. It is found that the interplay between spin-orbit coupling and repulsive interactions between anti-aligned pseudo-spins increases the population of the unaligned spin components in the ground state. The emergence of the mixed spin symmetric states compensates for the rise of the interaction energy. With the aligned interaction on, the avoided crossing between the ground state and the first excited state is observed only for small interaction, and this causes shape changes in the spin populations. Furthermore, we find that the pair correlation of the ground state shows similarly to that of Tonks-Girardeau gas even for relatively small contact interactions and such strong interaction feature is enhanced by the spin-orbit coupling.
@@ I Introduction
SOC in cold atoms is often discussed in the context of the mean-field regime, e.g. [8, 9, 10]. While the mean-field approximation has revealed interesting results, for example discovering the phase diagram [8], it imposes classical fields and ignores the quantum effects. To bridge single-particle physics and many-body physics in the SOC system, a mapping between the SOC cold atomic system and the Dicke model has been proposed [11]. However, the validity of the mapping is not obvious. The assumption imposed for the mapping is that all particles occupy the same real space state, which automatically leads to the pseudo-spin state spanned in only symmetric spin space. This assumption is commonly used for Bose gases with two internal degrees of freedom [12], however SOC couples pseudo-spin states and real space states and allows pseudo-spin states to get out of the symmetric spin space. It is not clear how even the lowest energy state is confined in symmetric spin space.
@@
The spin symmetry of the ground state in the two-particle system with SOC has been investigated by performing exact diagonalization [13]. Competition between SOC and contact interaction has been studied, and the emergence of anti-symmetric pseudo-spin states for strong interactions has been found. Despite using a bosonic system, anti-symmetric pseudo-spin states are preferable for the ground state. This work has also revealed the parameter regime where anti-symmetric pseudo-spin states or only symmetric pseudo-spin states are contained. However, these results cannot be extended to many-particle systems straightforwardly. One of the reasons is because two-pseudo-spin states have unique spin symmetry. Particles with two internal degrees of freedom are classified into symmetric and anti-symmetric pseudo-spin states, but the spin symmetry can be mixed in more-particle systems and anti-symmetry under all permutation of pseudo-spins no longer exist. It is unclear how the mixed spin symmetry affects the property of the ground state.
@@
In this work, we investigate the role of mixed spin symmetry in the SOC system. To this end, we consider a three-particle system with two internal states as it is the smallest system that has mixed spin symmetry and focus on comparison with a two-particle system. We have built the Hamiltonian of three bosons with pseudo-spin-1/2 in the presence of SOC by considering second quantisation and computed the ground state. Our numerical results confirm the appearance of mixed symmetric pseudo-spin states in the ground state for strong interaction between anti-aligned pseudo-spins and reveal the parameter regime where it is observed. Furthermore, the emergence of the mixed spin symmetry states reduces the interaction energy. This is also seen in the two-particle system when the anti-symmetric pseudo-spin states emerge. This implies that the mixed spin symmetric state mimics the anti-symmetric pseudo-spin state to suppress the energy. In addition, we study the spatial structure of the ground state by looking at the pair correlation and find that SOC assists contact interaction in inducing a strong interaction effect. Even for relatively small contact interaction, the pair correlation of the ground state has similarly to the strongly correlated Tonks-Girardeau gas. This is not observed in the two-particle system.
@@
The structure of this paper is as follows. Section II explains the basis of our system. First, we introduce the Hamiltonian of our system in the first quantization representation and then transform it into the second quantization representation. This is necessary to implement exact diagonalization. Next, we explain the spin symmetry in pseudo-spin-1/2 systems and clarify the existence of non-symmetric pseudo-spin states, which are often ignored. Section III shows our numerical results of the ground state and the energy spectrum. We consider two cases, (i) repulsive interaction between anti-aligned pseudo-spins and (ii) repulsive interaction between aligned pseudo-spins. The conclusion is given in Sec. IV.
II Few interacting bosons with SOC
@@
We present our formalism here. First, we introduce the Hamiltonian of our system in the first quantization basis and then rewrite it in the second quantization basis to induce the bosonic symmetry. This approach is often used to present the Hamiltonian of few particle systems, e.g. [14, 15, 16], and we refer interested readers to a detailed reference [17]. Second, we explain the spin symmetry of two spins and more spins respectively. While bosonic systems with two hyper-fine states are well studied, usually only pseudo-spin symmetric states are focused on, and the existence of anti-symmetry and mixed spin symmetry does not attract attention. We review the categories of spin symmetry before showing our numerical results.
II.1 Hamiltonian in the first quantisation basis
@@
$\hat{H}_{\downarrow\uparrow}$ $=\sum_{i<j}g_{\uparrow\uparrow}\delta(x_{i}-x_{j})$
@@
II.2 Hamiltonian in the second quantization basis
The above formalism is called the first quantization and describes distinguishable particles. To express indistinguishable particles simply, we implement the second quantization representation in a truncated space. Consider that $N$ particles occupy $N$ of first $M$ eigenstates of the harmonic oscillator for down spins and up spins. Defining the creation and annihilation operators for the $j$th eigenstate as $\hat{a}_{j}$ and $\hat{a}_{j}^{\dagger}$, the second quantised version of the single-particle part (2) is given by
@@ II.2 Hamiltonian in the second quantization basis
II.3 Spin symmetry of bosons with two internal states
Here, we review the symmetry of pseudo-spin-1/2 systems, and let us start with two-particle states. Since we consider bosons, every state is symmetric under permutation. The considered system has the pseudo-spin degree and the spatial degree of freedom, and the pseudo-spin basis are classified into three symmetric pseudo-spin states, $\ket{\downarrow\downarrow}$, (|+|$(\ket{\downarrow\uparrow}+\ket{\uparrow\downarrow})/\sqrt{2}$, $\ket{\uparrow\uparrow}$ and one anti-symmetric pseudo-spin state (||$(\ket{\downarrow\uparrow}-\ket{\uparrow\downarrow})/\sqrt{2}$. To satisfy the bosonic symmetry, the spatial states of symmetric pseudo-spin states are symmetric, i.e. even functions in the relative coordinate $x_{1}-x_{2}$. Similarly, the spatial states of anti-symmetric pseudo-spin states are anti-symmetric, i.e. odd functions in the relative coordinate $x_{1}-x_{2}$. Since the contact interaction occurs at a point and is described as the delta function $\delta(x_{1}-x_{2})$, the anti-symmetric spatial states do not feel the contact interaction.
|$\ket{S_{1}}$ $=\ket{\downarrow\downarrow\downarrow}$
|$\ket{S_{4}}$ $=\ket{\uparrow\uparrow\uparrow},$ (11)
and the following mixed spin symmetric basis,
For example, switching the first spin and the second spin in the pseudo-spin basis |$\ket{M_{1}}$ changes nothing (symmetric), but by switching the second spin and the third spin, the resultant state cannot be described with |$\ket{M_{1}}$ anymore. Therefore, these basis do not belong to symmetry or anti-symmetry but are called mixed symmetry. To satisfy symmetry under all permutation, the spatial states of the mixed spin symmetric states are neither symmetric nor anti-symmetric but formed such that the total state is symmetric.
When the spatial states and the pseudo-spin states are decoupled in systems of pseudo-spin-1/2 bosons, the pseudo-spin space is confined to the symmetric spin space. For instance, a pseudo-spin-1/2 BEC in a harmonic trap can be modelled with the collective spin operators, and the pseudo-spin Hamiltonian is expressed as $\hat{H}=g\hat{S}_{z}^{2}$ [12]. The influence of the spatial state is included in the interaction strength $g$, and the strength $g$ can be tuned by modifying the potential for example. However, SOC couples the spatial degree and the pseudo-spin degree, and the pseudo-spins can be non-symmetric. For instance, in the two-particle system, relative motion is induced by coupling between symmetric and anti-symmetric spin states (see Eq. (3) in Ref. [13]). Non-symmetric spin states play as an important role as symmetric spin states in the system with SOC.
III Numerical results
We construct the Hamiltonains (7)(II.2) and diagonalize them in a truncated space numerically. In this work, we fix $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}=g$ to keep the symmetry between down and up spins. Moreover, to study competition between SOC and contact interaction, we turn on the anti-aligned pseudo-spin interaction $g_{\downarrow\uparrow}$ or the aligned pseudo-spin interaction $g$. The study of the two particle system has shown that these two cases give clear different properties in the ground state [13]. We use the trap energy $\hbar\omega$ as the energy unit and normalise the parameters with the trap energy $\hbar\omega$ and the trap length $\xi=\sqrt{\hbar/m\omega}$, and display normalised parameters in all the figures. Also, we fix $k\xi=4$ and set the cutoff of the truncated space as $M=50$. We discuss the justification of the cutoff $M$ in Appendix A, and our numerical code can be viewed in Ref. [18].
@@ III Numerical results
III.1 Zero interactions $(g=g_{\downarrow\uparrow}=0)$
@@
We plot the energy difference $E_{j}-E_{0}$ between some excited states and the ground state for the noninteracting case ($g=g_{\downarrow\uparrow}=0$) for changing the coherent coupling strength $\Omega$ and reconfirm that the degeneracy in the ground state is resolved off from $\Omega_{\text{c}}^{0}$ (see the dashed line in Fig. 1(a)). Therefore, we define another critical point $\Omega_{\text{c}}$ for the trap system as a point where the energy difference $(E_{1}-E_{0})/\hbar\omega$ between the ground state and the first excited state is larger than $10^{-2}$ (see the dotted line). In the absence of coherent coupling, the ground states are four-fold degenerate, and their pseudo-spin states are aligned pseudo-spins such as (i) three down-spins and (ii) three up-spins and upaligned pseudo-spins such as (iii) two up-spins and one down-spin (iv) one up-spin and two down-spins. Positive (negative) momentum is induced to down (up) spins due to SOC. The coherent coupling mixes these states, and in the strong coherent coupling limit an equally-weighted superposition of all pseudo-spin states obtains the lowest energy.
@@
III.2 Anti-aligned interactions ($g_{\downarrow\uparrow}>0$ and $g=0$)
@@
III.2.1 Pseudo-spin population
@@
Furthermore, we have computed the population $p_{\text{M}1},p_{\text{M}2},p_{\text{M}3},p_{\text{M}4}$ of the mixed spin symmetric basis (II.3) (see the details of the calculation to Appendix B). Since these populations are the same due to our setting $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}$, we have plotted the sum $p_{\text{M}}=p_{\text{M}1}+p_{\text{M}2}+p_{\text{M}3}+p_{\text{M}4}$ (see the dotted lines in Fig. 2(a,b)). We have found that a large amount of the mixed spin symmetric states appear in the ground state for the strong interaction case (b) in Fig. 2. Considering $p_{\text{unaligned}}=p_{\text{S}2+\text{S}3}+p_{\text{M}}$ with $p_{\text{S}2+\text{S}3}$ the population of |$\ket{S_{2}}$ and |$\ket{S_{3}}$, this leads to the non-monotonic growth of $p_{\text{aligned}}$, $p_{\text{unaligned}}$. This is also seen in the two-particle system, and in that case it is originated from the emergence of the anti-symmetric pseudo-spin states [13]. The interaction energy increases for increasing the contact interaction strength in general, but the anti-symmetric pseudo-spin states do not feel contact interactions. Thus, the appearance of the anti-symmetric pseudo-spin states for strong contact interactions suppresses the interaction energy. Therefore, it is a question whether the mixed spin symmetric states also reduce the interaction energy in the case we consider, and the answer is that they do. We have calculated the interaction energy $E_{\text{int}}=\langle\hat{H}_{\text{int}}\rangle$ and plotted it as a function of $\Omega$, which displays a dent for strong interaction (see Fig. 2(c)). Finally, we have inspected the non-monotonic growth of $p_{\text{unaligned}}$ for a wide range of interaction $g_{\downarrow\uparrow}$ (see Fig. 3) and confirmed that the excess over $3/4$ is larger for stronger $g_{\downarrow\uparrow}$.

### v2

# Enhanced strong interaction effect in synthetic spin-orbit coupling with mixed spin symmetry

@@ Abstract
Synthetic spin-orbit coupling in cold atoms couples the pseudo-spin and spatial degrees of freedom, and therefore the inherent spin symmetry of the system plays an important role. In systems of two pseudo-spin degrees, two particles contain symmetric states and anti-symmetric states, but the spin symmetry can be mixed for more particles. This mixed spin symmetry has been overlooked and has not been investigated thoroughly. We study the role of mixed spin symmetry in the presence of spin-orbit coupling and consider the system of three bosons with two hyper-fine states trapped in a harmonic potential. We investigate the ground state and the energy spectrum by implementing exact diagonalization. Similarly to two-particle systems, the interplay between spin-orbit coupling and repulsive interactions between anti-aligned pseudo-spins increases the population of the unaligned spin components in the ground state. Also, the emergence of the mixed spin symmetric states compensates for the rise of the interaction energy. In contrast to two-particle systems, the pair correlation of the ground state is analogous to the Tonks-Girardeau gas even for relatively small contact interactions, and this feature is enhanced by the spin-orbit coupling.
@@ I Introduction
SOC in cold atoms is often discussed in the context of the mean-field regime, e.g. [8, 9, 10]. While the mean-field approximation has revealed interesting results, for example discovering the phase diagram [8], it imposes classical fields and ignores the quantum effects. To bridge single-particle physics and many-body physics in SOC systems, a mapping between cold atomic SOC systems and the Dicke model has been proposed [11]. However, the validity of the mapping is not obvious [12]. The assumption imposed for the mapping is that all particles occupy the same real space state, which automatically leads to the pseudo-spin state spanning symmetric spin space only. This assumption is commonly used for Bose gases with two internal degrees of freedom [13], however SOC couples pseudo-spin states and real space states and allows pseudo-spin states to get out of symmetric spin space. It is not clear how the lowest energy state is confined in symmetric spin space.
@@
Motivated by this contradiction, we tackle the problem of multi-particle systems with SOC by computing low energy states exactly. The two-particle system with SOC in one dimension has already been investigated by performing exact diagonalization, and the competition between SOC and contact interactions has also been studied [14]. It has been shown that the ground state includes the anti-symmetric pseudo-spin state, which does not feel contact interactions. However, these results cannot be extended to many-particle systems straightforwardly, because they are rooted in the spin symmetry of only two particles. While two pseudo-spin states are classified into either symmetric or anti-symmetric pseudo-spin states, the spin symmetry can be mixed in more-particle systems, and it is unclear whether the properties of the ground state in the two-particle system are inherited in more-particle systems. Moreover, pseudo-spin states that are not symmetric have not had as much attention as symmetric pseudo-spin states and have not been investigated as thoroughly.
@@
In this work, we investigate the role of mixed spin symmetry in SOC systems. To this end, we consider the smallest system with mixed spin symmetry, a three-particle system with two internal states, and we focus on the comparison with two-particle systems. We have built the Hamiltonian of three bosons with two internal states in the presence of SOC by considering the second quantisation and have computed a few low-energy states. In comparison to two particle systems, we have found similarity in the appearance of mixed symmetric pseudo-spin states in the ground state, which is confirmed for strong interaction between anti-aligned pseudo-spins and reduced interaction energy. This is also seen in two-particle systems when the anti-symmetric pseudo-spin states emerge. Also, the emergence of non-symmetric pseudo-spin states challenges the mapping proposed in Ref. [11]. At the same time, we have also studied the spatial structure of the ground state by looking at the pair correlation and find that SOC assists contact interaction in inducing a strong interaction effect. Even for relatively small contact interactions, the pair correlation of the ground state has a structure similar to the strongly correlated Tonks-Girardeau gas. The pair correlation of the two-particle system is simply the total density distribution and does not display such a strong interaction effect.
@@
The structure of this paper is as follows. First, in Sec. II we review the spin symmetry in pseudo-spin-1/2 systems and clarify the existence of non-symmetric pseudo-spin states. Section III explains the formulation of our system. We introduce the Hamiltonian of our system in the first quantization representation and then transform it into the second quantization representation. This is necessary to implement exact diagonalization. Section IV shows our numerical results of the ground state and the energy spectrum. We consider two cases, (i) repulsive interaction between anti-aligned pseudo-spins and (ii) repulsive interaction between aligned pseudo-spins. The conclusion is given in Sec. V.
II Spin symmetry of bosons with two internal states
Before the main part of this work, we will review the symmetry of pseudo-spin-1/2 systems and clarify what mixed spin symmetry means in our context. First, let us start with two-particle states. Since we are considering bosons, every state is symmetric under permutation. The considered system has pseudo-spin and spatial degrees of freedom, and the pseudo-spin basis are classified into three symmetric pseudo-spin states, $\ket{\downarrow\downarrow}$, (|+|$(\ket{\downarrow\uparrow}+\ket{\uparrow\downarrow})/\sqrt{2}$, $\ket{\uparrow\uparrow}$ and one anti-symmetric pseudo-spin state (||$(\ket{\downarrow\uparrow}-\ket{\uparrow\downarrow})/\sqrt{2}$. To satisfy the bosonic symmetry, the spatial states of symmetric pseudo-spin states are symmetric, i.e. even functions in the relative coordinate $x_{1}-x_{2}$. Similarly, the spatial states of anti-symmetric pseudo-spin states are anti-symmetric, i.e. odd functions in the relative coordinate $x_{1}-x_{2}$. Since the contact interaction occurs at a point and is described as the delta function $\delta(x_{1}-x_{2})$, the anti-symmetric spatial states do not feel the contact interaction.
|$\ket{S_{1}}$ $=\ket{\downarrow\downarrow\downarrow}$
|$\ket{S_{4}}$ $=\ket{\uparrow\uparrow\uparrow},$ (1)
and the following mixed spin symmetric basis,
For example, switching the first spin and the second spin in the pseudo-spin basis |$\ket{M_{1}}$ changes nothing (symmetric), but by switching the second spin and the third spin, the resultant state can no longer be described with |$\ket{M_{1}}$. Therefore, these bases are not symmetric or anti-symmetric, but are instead mixed symmetric. To satisfy symmetry under all permutations, the spatial states of the mixed spin symmetric states are formed such that the total state is symmetric.
When spatial states and the pseudo-spin states are decoupled in systems of pseudo-spin-1/2 bosons, the pseudo-spin space is confined to the symmetric spin space. For instance, a pseudo-spin-1/2 Bose-Einstein condensate in a harmonic trap can be modelled with the collective spin operators, and the pseudo-spin Hamiltonian is expressed as $\hat{H}=g\hat{S}_{z}^{2}$ [13]. The influence of the spatial state is included in the interaction strength $g$, and the strength $g$ can be tuned by modifying the potential for example. However, as we introduce in the next section, SOC couples the spatial degree and the pseudo-spin degree. Thus, non-symmetric spin states play as an important role as symmetric spin states in the system with SOC.
III Few interacting bosons with SOC
@@
We present our formalism here. First, we introduce the Hamiltonian of our system in the first quantization basis and then rewrite it in the second quantization basis to induce the bosonic symmetry. This approach is often used to present the Hamiltonian of few particle systems, e.g. [15, 16, 17], and we refer interested readers to a detailed reference [18].
III.1 Hamiltonian in the first quantisation basis
@@
$\hat{H}_{\downarrow\uparrow}$ $=\sum_{i<j}g_{\downarrow\uparrow}\delta(x_{i}-x_{j})$
@@
The above formalism is called the first quantization and labels each particle as seen in Eqs. (4)(6). This is suitable to distinguishable particles although it can work for indistinguishable particles after some modification, for example see Ref. [14]. To express indistinguishable particles simply, we implement the second quantization representation in a truncated space.
III.2 Hamiltonian in the second quantization basis
Consider $N$ particles occupying $N$ of first $M$ eigenstates of the harmonic oscillator for down spins and up spins. Defining the creation and annihilation operators for the $j$th eigenstate as $\hat{a}_{j}$ and $\hat{a}_{j}^{\dagger}$, the second quantised version of the single-particle part (4) is given by
@@ III.2 Hamiltonian in the second quantization basis
IV Numerical results
We construct the Hamiltonains (7)(III.2) and diagonalize them in a truncated space numerically. In this work, we fix $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}=g$ to keep the symmetry between down and up spins. Moreover, to study competition between SOC and contact interaction, we turn on the anti-aligned pseudo-spin interaction $g_{\downarrow\uparrow}$ or the aligned pseudo-spin interaction $g$. The study of the two particle system has shown that these two cases give clear different properties in the ground state [14]. We use the trap energy $\hbar\omega$ as the energy unit and normalise the parameters with the trap energy $\hbar\omega$ and the trap length $\xi=\sqrt{\hbar/m\omega}$, and display normalised parameters in all the figures. Also, we fix $k\xi=4$ and set the cutoff of the truncated space as $M=50$. We discuss the justification of the cutoff $M$ in Appendix A, and our numerical code can be viewed in Ref. [19].
@@ IV Numerical results
IV.1 Zero interactions $(g=g_{\downarrow\uparrow}=0)$
@@
We plot the energy difference $E_{j}-E_{0}$ between some excited states and the ground state for the noninteracting case ($g=g_{\downarrow\uparrow}=0$) for changing the coherent coupling strength $\Omega$ and reconfirm that the degeneracy in the ground state is resolved off from $\Omega_{\text{c}}^{0}$ (see the dashed line in Fig. 1(a)). Therefore, we define another critical point $\Omega_{\text{c}}$ for the trap system as a point where the energy difference $(E_{1}-E_{0})/\hbar\omega$ between the ground state and the first excited state is larger than $10^{-2}$ (see the dotted line). In the absence of coherent coupling, the ground states are four-fold degenerate, and their pseudo-spin states are aligned pseudo-spins such as (i) $\ket{\uparrow\uparrow\uparrow}$ and (ii) $\ket{\downarrow\downarrow\downarrow}$ and unaligned pseudo-spins such as (iii) $\ket{\uparrow\uparrow\downarrow}+\ket{\uparrow\downarrow\uparrow}+\ket{\downarrow\uparrow\uparrow}$ and (iv) $\ket{\uparrow\downarrow\downarrow}+\ket{\downarrow\uparrow\downarrow}+\ket{\downarrow\downarrow\uparrow}$. Positive (negative) momentum is induced to down (up) spins due to SOC. The coherent coupling mixes these states, and in the strong coherent coupling limit an equally-weighted superposition of all pseudo-spin states obtains the lowest energy.
@@
IV.2 Anti-aligned interactions ($g_{\downarrow\uparrow}>0$ and $g=0$)
@@
IV.2.1 Pseudo-spin population
@@
Furthermore, we have computed the population $p_{\text{M}1},p_{\text{M}2},p_{\text{M}3},p_{\text{M}4}$ of the mixed spin symmetric basis (II) (see the details of the calculation to Appendix B). Since these populations are the same due to our setting $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}$, we have plotted the sum $p_{\text{M}}=p_{\text{M}1}+p_{\text{M}2}+p_{\text{M}3}+p_{\text{M}4}$ (see the dotted lines in Fig. 2(a,b)). We have found that a large amount of the mixed spin symmetric states appear in the ground state for the strong interaction case (b) in Fig. 2. Considering $p_{\text{unaligned}}=p_{\text{S}2+\text{S}3}+p_{\text{M}}$ with $p_{\text{S}2+\text{S}3}$ the population of |$\ket{S_{2}}$ and |$\ket{S_{3}}$, this leads to the non-monotonic growth of $p_{\text{aligned}}$, $p_{\text{unaligned}}$. This is also seen in the two-particle system, and in that case it is originated from the emergence of the anti-symmetric pseudo-spin states [14]. The interaction energy increases for increasing the contact interaction strength in general, but the anti-symmetric pseudo-spin states do not feel contact interactions. Thus, the appearance of the anti-symmetric pseudo-spin states for strong contact interactions suppresses the interaction energy. In three-particle systems, the mixed spin symmetric states is mixture of spin symmetric states and spin antisymmetric states, and this anti-symmetry reduces the interaction energy. We have calculated the interaction energy $E_{\text{int}}=\langle\hat{H}_{\text{int}}\rangle$ and plotted it as a function of $\Omega$, which displays a dent for strong interaction (see Fig. 2(c)). Finally, we have inspected the non-monotonic growth of $p_{\text{unaligned}}$ for a wide range of interaction $g_{\downarrow\uparrow}$ (see Fig. 3) and confirmed that the excess over $3/4$ is larger for stronger $g_{\downarrow\uparrow}$.